OpenStudy (anonymous):

Help me please-->> Solve: 3(x + 1) + 6 = 33

OpenStudy (anonymous):

@AriPotta @Argonx16 @AccessDenied @andreadesirepen @Alyx_da_Boss @satellite73 @drummer07 Anyone plz help me

OpenStudy (aripotta):

hey, i already did that lol. look on the other thread

OpenStudy (anonymous):

Oh i got it answers 8 but who want the medal??

OpenStudy (anonymous):

Step 1: 3(x + 1) + 6 = 33 Distribute 3 over to x+1 You get 3x+3+6=33 add the 6 and 3 3x+9=33 Subtract the 9

OpenStudy (anonymous):

And i need help with the next question anyways